# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ A.2 Proof of Lemma 14
$g(t)=\bm{M}-\prod_{i=1}^{L}(\bm{W}_{i}+t\bm{U}_{i}),\quad f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})=\Bigl<g(t),g(t)\Bigr$>. (49)
First, we need to differentiate $f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})$ w.r.t $t$. Using the fact that $\Bigl<\bm{A},\bm{B}\Bigr>=\operatorname{tr}(\bm{A}^{\top}\bm{B})$, which simplifies the differentiation,
$\frac{d}{dt}f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})=2\Bigl<g(t),g^{\prime}(t)\Bigr$>. (50)
$\frac{d^{2}}{dt^{2}}f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})=2\Bigl<g^{\prime}(t),g^{\prime}(t)\Bigr>+2\Bigl<g(t),g^{\prime\prime}(t)\Bigr$>. (51)
Then, the directional second derivative $\nabla^{2}f(\bm{W}_{1},\cdots,\bm{W}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]$ equals
$\frac{d^{2}}{dt^{2}}f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})\Bigr|_{t=0}=2\Bigl<g^{\prime}(0),g^{\prime}(0)\Bigr>+2\Bigl<g(0),g^{\prime\prime}(0)\Bigr$>. (52)
@@
$g(0)$ $=\bm{M}-\prod_{i=1}^{L}\bm{W}_{i},$ (53)
$g^{\prime}(0)$ $=-\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}\Bigr)\Biggr],$ (54)
$g^{\prime\prime}(0)$ $=-2\sum_{1\leq k<i\leq L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=k+1}^{i-1}\bm{W}_{j}\Bigr)\bm{U}_{k}\Bigl(\prod_{j=1}^{k-1}\bm{W}_{j}\Bigr)\Biggr].$ (55)
Therefore, for any $\bm{W}_{1},\bm{W}_{2},\cdots,\bm{W}_{L}]$ in parameter space
$\nabla^{2}f(\bm{W}_{1},\cdots,\bm{W}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]\hfill=$ (56)
$2\Bigl<\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}\Bigr)\Biggr],\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}\Bigr)\Biggr]\Bigr$> (57)
- $4\Bigl<\bm{M}-\prod_{i=1}^{L}\bm{W}_{i},\sum_{1\leq k<i\leq L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=k+1}^{i-1}\bm{W}_{j}\Bigr)\bm{U}_{k}\Bigl(\prod_{j=1}^{k-1}\bm{W}_{j}\Bigr)\Biggr]\Bigr$>. (58)
Note that for any minimizer $\bm{W}^{*}_{1},\bm{W}^{*}_{2},\cdots,\bm{W}^{*}_{L}]$, $\bm{M}-\prod_{j=1}^{L}\bm{W}^{*}_{i}=0$. Hence, for any global minimum
$\nabla^{2}f(\bm{W}^{*}_{1},\cdots,\bm{W}^{*}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]=\hfill$ (59)
$2\Bigl<\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr],\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}^{*}_{j}\Bigr)\Biggr]\Bigr$>. (60)
@@ Appendix B Warm-Up: Deep Overparameterized Scalar Factorization
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)$ $=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (61)
@@
$\leq\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\dots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\norm{\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)}_{F}\Biggr)^{2}$ (62)
@@
$=\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\norm{\Bigl[\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\Bigr]\bm{u}_{i}}_{2}\Biggr)^{2}$ (63)
@@
$\leq\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Biggl(\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\Biggr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2}$ (64)
@@
$=\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2},$ (65)
@@
where $\bm{U}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $\bm{u}_{i}\in\mathbb{R}^{d_{i}d_{i-1}}$. We can upper bound the right-hand side of (61) by using the triangle inequality. By applying the vectorization trick of the Kronecker product, we can rewrite (62). Then, noting the fact that for any matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and vector $\bm{x}\in\mathbb{R}^{n}$, $\norm{\bm{Ax}}_{2}\leq\sigma_{\max}(\bm{A})\norm{\bm{x}}_{2}$, we can upper bound the right-hand side of (63). Note that for any matrix $\bm{A}$ and $\bm{B}$, $\sigma_{\max}(\bm{A}\otimes\bm{B})=\sigma_{\max}(\bm{A})\sigma_{\max}(\bm{B})$. Hence, we can rewrite (64). Then, by using the Cauchy–Schwarz inequality,
@@
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)\leq 2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2}.$ (66)
@@
Then, it suffices to show that there exists a direction $\bm{U}_{1}^{*},\bm{U}_{2}^{*},\dots,\bm{U}_{L}^{*}]$ along which the bound in (66) is achieved.
@@
Consider decomposition of $\prod_{j=i+1}^{L}\bm{W}^{*}_{j}$ by SVD, and denote by $\bm{u}_{L_{i}}$ and $\bm{v}_{L_{i}}$ the left and right singular vectors of $\prod_{j=i+1}^{L}\bm{W}^{*}_{j}$ corresponding to the largest singular value, respectively. Note that since $\bm{W}_{L}$ is a vector, we have $\bm{u}_{L_{i}}=1$ for all $i\in[L]$. Moreover, decompose $\prod_{j=1}^{i-1}\bm{W}_{j}^{*}$ by SVD, and denote by $\bm{u}_{R_{i}}$ and $\bm{v}_{R_{i}}$ the left and right singular vectors of $\prod_{j=1}^{i-1}\bm{W}_{j}^{*}$ corresponding to the largest singular value, respectively. Note that since $\bm{W}_{1}$ is a vector, we have $\bm{v}_{R_{i}}=1$ for all $i\in[L]$. Now, we determine a particular direction $\bm{U}_{1}^{*},\bm{U}_{2}^{*},\dots,\bm{U}_{L}^{*}]$ such that they achieve the upper bound while satisfying the constraint $\sum_{i=1}^{L}\norm{\bm{U}_{i}^{*}}_{F}^{2}=1$. Choose
$\bm{U}_{i}^{*}=\frac{\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)}{\sqrt{\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2}}}\bm{v}_{L_{i}}\bm{u}_{R_{i}}^{\top}.$ (67)
@@
$2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}^{*}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2}.$ (68)
@@
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w^{*}}))=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2}.$ (69)

### v5

@@ A.1 Proof of Lemma 14
$g(t)=\mathbf{M}-\prod_{i=1}^{L}(\mathbf{W}_{i}+t\mathbf{U}_{i}),\quad f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})=\Bigl<g(t),g(t)\Bigr$>. (52)
First, we need to differentiate $f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})$ with respect to $t$. Using the fact that $\Bigl<\mathbf{A},\mathbf{B}\Bigr>=\operatorname{tr}(\mathbf{A}^{\top}\mathbf{B})$, which simplifies the differentiation,
$\frac{d}{dt}f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})=2\Bigl<g(t),g^{\prime}(t)\Bigr$>. (53)
$\frac{d^{2}}{dt^{2}}f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})=2\Bigl<g^{\prime}(t),g^{\prime}(t)\Bigr>+2\Bigl<g(t),g^{\prime\prime}(t)\Bigr$>. (54)
Then, the directional second derivative $\nabla^{2}f(\mathbf{W}_{1},\cdots,\mathbf{W}_{L})[\mathbf{U}_{1},\cdots,\mathbf{U}_{L}]$ equals
$\frac{d^{2}}{dt^{2}}f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})\Bigr|_{t=0}=2\Bigl<g^{\prime}(0),g^{\prime}(0)\Bigr>+2\Bigl<g(0),g^{\prime\prime}(0)\Bigr$>. (55)
@@
$g(0)$ $=\mathbf{M}-\prod_{i=1}^{L}\mathbf{W}_{i},$ (56)
$g^{\prime}(0)$ $=-\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}\Bigr)\Biggr],$ (57)
$g^{\prime\prime}(0)$ $=-2\sum_{1\leq k<i\leq L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=k+1}^{i-1}\mathbf{W}_{j}\Bigr)\mathbf{U}_{k}\Bigl(\prod_{j=1}^{k-1}\mathbf{W}_{j}\Bigr)\Biggr].$ (58)
Therefore, for any $\mathbf{W}_{1},\mathbf{W}_{2},\cdots,\mathbf{W}_{L}]$ in parameter space
$\nabla^{2}f(\mathbf{W}_{1},\cdots,\mathbf{W}_{L})[\mathbf{U}_{1},\cdots,\mathbf{U}_{L}]\hfill=$ (59)
$2\Bigl<\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}\Bigr)\Biggr],\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}\Bigr)\Biggr]\Bigr$> (60)
- $4\Bigl<\mathbf{M}-\prod_{i=1}^{L}\mathbf{W}_{i},\sum_{1\leq k<i\leq L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=k+1}^{i-1}\mathbf{W}_{j}\Bigr)\mathbf{U}_{k}\Bigl(\prod_{j=1}^{k-1}\mathbf{W}_{j}\Bigr)\Biggr]\Bigr$>. (61)
Note that for any minimizer $\mathbf{W}^{*}_{1},\mathbf{W}^{*}_{2},\cdots,\mathbf{W}^{*}_{L}]$, $\mathbf{M}-\prod_{j=1}^{L}\mathbf{W}^{*}_{i}=0$. Hence, for any global minimum
$\nabla^{2}f(\mathbf{W}^{*}_{1},\cdots,\mathbf{W}^{*}_{L})[\mathbf{U}_{1},\cdots,\mathbf{U}_{L}]=\hfill$ (62)
$2\Bigl<\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr],\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}^{*}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}^{*}_{j}\Bigr)\Biggr]\Bigr$>. (63)
@@ Appendix B Warm-Up: Deep Overparameterized Scalar Factorization
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)$ $=\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}$ (64)
@@
$\leq\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\dots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\norm{\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)}_{F}\Biggr)^{2}$ (65)
@@
$=\max_{\begin{subarray}{c}\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\norm{\Bigl[\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\Bigr]\mathbf{u}_{i}}_{2}\Biggr)^{2}$ (66)
@@
$\leq\max_{\begin{subarray}{c}\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Biggl(\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{\top}\otimes\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\Biggr)\norm{\mathbf{u}_{i}}_{2}\Biggr)^{2}$ (67)
@@
$=\max_{\begin{subarray}{c}\mathbf{u}_{1},\mathbf{u}_{2},\dots,\mathbf{u}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\norm{\mathbf{u}_{i}}_{2}\Biggr)^{2},$ (68)
@@
where $\mathbf{U}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $\mathbf{u}_{i}\in\mathbb{R}^{d_{i}d_{i-1}}$. We can upper bound the right-hand side of (64) by using the triangle inequality. By applying the vectorization trick of the Kronecker product, we can rewrite (65). Then, noting the fact that for any matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ and vector $\mathbf{x}\in\mathbb{R}^{n}$, $\norm{\mathbf{Ax}}_{2}\leq\sigma_{\max}(\mathbf{A})\norm{\mathbf{x}}_{2}$, we can upper bound the right-hand side of (66). Note that for any matrix $\mathbf{A}$ and $\mathbf{B}$, $\sigma_{\max}(\mathbf{A}\otimes\mathbf{B})=\sigma_{\max}(\mathbf{A})\sigma_{\max}(\mathbf{B})$. Hence, we can rewrite (67). Then, by using the Cauchy–Schwarz inequality,
@@
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)\leq 2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{2}.$ (69)
@@
Then, it suffices to show that there exists a direction $\mathbf{U}_{1}^{*},\mathbf{U}_{2}^{*},\dots,\mathbf{U}_{L}^{*}]$ along which the bound in (69) is achieved.
@@
Consider decomposition of $\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}$ by SVD, and denote by $\mathbf{u}_{L_{i}}$ and $\mathbf{v}_{L_{i}}$ the left and right singular vectors of $\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}$ corresponding to the largest singular value, respectively. Note that since $\mathbf{W}_{L}$ is a vector, we have $\mathbf{u}_{L_{i}}=1$ for all $i\in[L]$. Moreover, decompose $\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}$ by SVD, and denote by $\mathbf{u}_{R_{i}}$ and $\mathbf{v}_{R_{i}}$ the left and right singular vectors of $\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}$ corresponding to the largest singular value, respectively. Note that since $\mathbf{W}_{1}$ is a vector, we have $\mathbf{v}_{R_{i}}=1$ for all $i\in[L]$. Now, we determine a particular direction $\mathbf{U}_{1}^{*},\mathbf{U}_{2}^{*},\dots,\mathbf{U}_{L}^{*}]$ such that they achieve the upper bound while satisfying the constraint $\sum_{i=1}^{L}\norm{\mathbf{U}_{i}^{*}}_{F}^{2}=1$. Choose
$\mathbf{U}_{i}^{*}=\frac{\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)}{\sqrt{\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{2}}}\mathbf{v}_{L_{i}}\mathbf{u}_{R_{i}}^{\top}.$ (70)
@@
$2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\Bigr)\mathbf{U}^{*}_{i}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{2}.$ (71)
@@
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\mathbf{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\Bigr)^{2}.$ (72)